The variables x and y vary directly. Use the given values to write an equation that relates x and y:
x = 4, y = -16

Answer by ewatrrr(24785)

 
Hi,
The variables x and y vary directly
y = kx
-16 = k·4 || when x = 4, y = -16
-4 = k
y = -4x